# Infinite-Horizon Optimal Wireless Control Over Shared State-Dependent Fading Channels for IIoT Systems

### v1

@@ 2.3 Multiloop WCS
The transmission of measurements and control decisions might fail due to packet decoding errors and packet collisions. In this paper, we focus on the uplink transmission from the sensors to the access point, the idea can be extended into the downlink transmission scenario straightforwardly. Use $\lambda_{i}(k)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time $k$. Then, the measurement $x_{i}(k)$ is used to compute the control input $v_{i}(k)$ by pre-designed control policy at a successful transmission ($\lambda_{i}(k)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$. Otherwise, an open-loop system with system matrix $A_{o,i}$ which may be unstable is obtained. The resulting system can be uniformly described as the following switched model:
$x_{i}(k+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=1,\\ A_{o,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
@@ 2.4 Finite-field MAS
Due to the discretization of the workspace, it is the cell that the agents are located in, not the specific position, that directly affects the channel condition of the WCS. In this way, the state of an agent is defined as the index of the cell that the agent is located in, i.e., the state space of each agent is the finite set $\mathcal{D}_{\kappa}$.
At each time step, each agent updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite field model:
$\alpha_{j}(k+1)=\{_{\kappa}\sum\limits_{l\in\bar{\mathcal{I}}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
where ¯$\bar{\mathcal{I}}_{j}=\mathcal{I}_{j}\cup\{j\}$, $\mathcal{I}_{j}$ is the set of in-neighbors of $j$, $\alpha_{j}(k)$, $u_{j}(k)$ respectively denote the state and input of $j$ at time $k$, $a_{j,l}\in\mathcal{D}_{\kappa}$ are the weights, and operations $+_{\kappa}$, ×$\times_{\kappa}$ denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$, respectively [Lidl1996]. In this way, the state $\alpha_{j}(k)$ will evolve over the finite cells $\{0,1,\cdots,\kappa-1\}$. Such finite field architecture has been widely studied to address problems for networks of agents with finite state space [Pasqualetti2014, Sundaram2013]. To ensure production safety such as collision and dangerous position avoidance, we impose state constraints $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}^{n}$ and state-dependent input constraints $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))\subseteq\mathcal{D}_{\kappa}^{n}$. This is one of the differences with the existing results [hu2022, hu2019] which model the mobile agent by an unconstrained Markov decision process. The existence of constraints makes the control of MAS much more complex.
@@ 2.5 State-dependent fading channel model
In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and thus influence the local channel state. We call a packet collision happens on system $i$, rendering system $i$’s packet decoding failed, if some other sensor $j\neq i$ transmits in the same time slot. If sensor $i$ transmits and has a collision-free time slot, the success of decoding the packet at the access point depends on the local channel state of link $i$, then the state of the MAS. Such WCS and MAS coupling is modeled by the following state-dependent fading channel:
$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\lambda}_{i}(\alpha),$ (5)
where $\bar{\lambda}_{i}(\alpha):\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$, $\lambda_{i}(k)=1$ indicates the successful transmission of system $i$.
Eq. (5) presents a general state-dependent channel model. Communication policies for sensors and the interference among transmission links can be embedded into this model. In this model, the probability of a successful transmission for WCS (3) depends on the state of MAS (4). It is quite different from the traditional shadow fading model [Gatsis2014, zhangq1999] assuming the channel states independent from the physical states. Such coupling makes the traditional methods on performance analysis of WCSs [gatsis2015, gatsis2018, Hristu2001, le2011, Molin, zhang2006] fail in the scenario in this paper.
@@ 2.6 Problem formulation
Considering the state-dependent fading channel model (5), given initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$ of the MAS, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ for MAS (4) such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k\in\mathbb{N}$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
@@
The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}\subseteq\mathcal{D}_{\kappa}^{n}$ and a positive integer $K$ such that the inequalities
$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]\leq\rho_{i}V_{i}(x_{i}(k))$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+Tr(Q_{i}\Xi_{i}),\forall\ x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ (6)
hold for any $k\geq K$, any $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ and any $i\in\{1,\cdots,q\}$, where $\mathbb{E}$ is the expectation, $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
In (6), $\rho_{i}$ represents desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation. Similar kinds of performance requirements as in (6) have been adopted for WCSs [gatsis2015, gatsis2018]. However, different from them, Definition 1 explicitly explores the state-dependent fading channel.
@@
$J_{{\bf{u}}}(\alpha_{0})=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\sum\limits_{k=0}^{K-1}\Big{\{}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda g(\alpha(k;\alpha_{0},{\bf{u}}),u(k))\Big{\}}$ (7)
reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}^{m+n}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},$
$~{}~{}~{}~{}~{}~{}~{}~{}u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (8)
In the above problem, there are heterogeneous system dynamics constraints (3), (4), WCS’s performance requirements (6) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, which make the resolution of this problem difficult.
Figure 2: Main idea of solving the infinite-horizon optimal control problem (8).
3 Lyapunov-like performance analysis over the state-dependent fading channel
In the optimal control problem (8), Lyapunov-like performance requirements constraints (6) brings about couplings between the WCS and the MAS, making the resolution of problem (8) challenging. To address this issue, we convert constraint (6) for WCS (3) and safety constraints to the constrained set stabilization of MAS (4) using the state-dependent fading channel (5). In this way, all constraints in problem (8) are related to the MAS, which facilitates its optimal input sequence design. Motivated by this observation, the main idea of solving problem (8) is illustrated in Fig. 2.
3.1 Characterization of state-dependent fading channel
In this part, we characterize the state-dependent fading channel in (5) by taking into account of the wireless communication policy and the interference among transmission links.
For fixed wireless communication policy, the probability of successful transmission for link $i$ can be expressed as $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}$, where $\eta_{i}(k)=1$ indicates the successful packet decoding. Due to the discrete channel state, the local channel state in the presence of shadow fading caused by movements of mobile agents is computed as follows:
$\mathbb{P}\{\gamma_{i}(k)=a|\alpha(k)=\alpha\}=\bar{\gamma}_{i}(a,\alpha),i=1,\cdots,q,$ (9)
where $\mathbb{P}$ is the probability, and $\bar{\gamma}_{i}:\mathcal{D}_{r}\times\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$ can be easily obtained by measuring the local fading amplitude for each MAS state according to the statistical methods in [Agrawal2014, Kashiwagi2010] and reference therein. Then, according to (2), we have $\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)$
Since the transmit power is fixed, the probability that the packet of system $i$ is successfully decoded at the access point depends on the local channel state of link $i$, then the state of the MAS as
$\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\eta}_{i}(\alpha),$ (10)
where $\bar{\eta}_{i}:\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$. The interference among transmission links is taken into account by adopting the collision channel model. Then, it holds
$\bar{\eta}_{i}(\alpha)=$$\mathbb{P}\{\hbox{SNR}_{i}(k)\geq\eta_{i}^{0}|\alpha(k)=\alpha,\varphi_{i}(k)=1,$
$\varphi_{j}(k)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(k)=0|\alpha(k)=\alpha\},$
where $\hbox{SNR}_{i}(k)=\mu_{i}\gamma_{i}^{2}(k)/N_{0}$, $N_{0}$ and $\eta_{i}^{0}$ respectively denote the signal-to-noise ratio [hu2019], noise power and the specified decodability threshold of system $i$. Based on the above analysis, we obtain the following result.
Lemma 2.
For the state-dependent fading channel (5), it holds $\bar{\lambda}_{i}(\alpha)=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)\bar{\eta}_{i}(\alpha)$, $\alpha\in\mathcal{C}_{\alpha}$, $i=1,\cdots,q$.
3.2 Lyapunov-like performance analysis of WCS
Based on the characterization of state-dependent fading channel, we can now analyze the Lyapunov-like performance of the WCS. For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ an admissible input sequence in the following.

### v2

@@ 2.3 Multiloop WCS
The transmission of measurements and control decisions might fail due to packet decoding errors and packet collisions. Assume that both wireless uplink and downlink have the same channel characteristics [hu2019]. In this case, we focus on the uplink transmission from the sensors to the access point, and the idea can be extended into the downlink transmission scenario straightforwardly. Use $\lambda_{i}(l)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time instant $t_{l}$. Then, the measurement $x_{i}(l)$ is used to compute the control input $v_{i}(l)$ at a successful transmission ($\lambda_{i}(l)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$. Otherwise, an open-loop system with the system matrix $A_{o,i}$, which may be unstable, is obtained. The resulting system can be described as the following switched model:
$x_{i}(l+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=1,\\ A_{o,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
@@ 2.4 Finite-field MAS
An MAS coordinates with the multiloop WCS to jointly perform smart manufacturing tasks. Mobile agents are driven by the standard single-integrator dynamics222 This paper uses the model as an illustration of how to abstract the finite-field MAS from the physics dynamics. The proposed results are general and can be extended to other models for mobile agents. as
$z_{j}(k+1)=z_{j}(k)+T\nu_{j}(k),j=1,\cdots,n,$
where $z_{j}(k),\nu_{j}(k)\in\mathbb{R}^{2}$ respectively denote the position and velocity of agent $j$ in the two-dimensional space at time instant $t_{\tau k}$, $\{t_{\tau k}\}_{k=0}^{\infty}$ is the sequence of sampling time instants with time step $T$, and $\tau$ is a positive integer indicating different time steps for the evolution of the MAS and WCS.
When coordinating with the WCS to jointly perform overall tasks, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over. Note that what matters for shadow fading is the position in space within certain intervals [Ahlen2019], not the specific position determined by the single-integrator dynamics. Partition the whole workspace into $\kappa$ non-overlapping cells $\{0,1,\cdots,\kappa-1\}$. Abstract the cell in which each mobile agent is located as a logical state in $\mathcal{D}_{\kappa}$. Each agent interacts with others by on-board sensors, and updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite-field model:
$\alpha_{j}(k+1)\!=\!\{_{\kappa}\!\sum\limits_{l\in\mathcal{I}_{j}\cup\{j\}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
where $\mathcal{I}_{j}$ is the set of all in-neighbors of $j$, $a_{j,l}\in\mathcal{D}_{\kappa}$ denotes the weight, operations $+_{\kappa}$ and ×$\times_{\kappa}$ respectively denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$ [Lidl1996], $\alpha_{j}(k)$, $u_{j}(k)\in\mathcal{D}_{\kappa}$ respectively denote the state and control input of agent $j$ at time instant $t_{\tau k}$. The feedback term $\{_{\kappa}\sum_{l\in\mathcal{I}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)$ together with $u_{j}(k)$ can describe several mobile agents’ controllers, including consensus, rendezvous and formation [Mesbahi, Pasqualetti2014, Sundaram2013]. The constraint on state $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}$ with $\alpha_{j}(k)\in\mathcal{C}_{\alpha_{j}}$, $j=1,\cdots,n$ is imposed to ensure each agent $j$ moves within its task areas $\mathcal{C}_{\alpha_{j}}$ while avoiding dangerous positions, such as collisions with industrial plants. In addition, the transition constraint $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))$ is imposed to restrict agents’ movements within the actual velocity limits. Such safety constraints distinguish the model from previous works [hu2022, hu2019] representing the mobile agent by an unconstrained Markov decision process.
@@ 2.5 State-dependent fading channel model
In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and finally influence the probability of successful wireless transmission. With different time steps, during each sampling interval $[t_{\tau k},t_{\tau(k+1)})$ of the MAS, there are $\tau$ time instants of the wireless transmission as $t_{l}$, $l={\tau k},{\tau k+1},\cdots,{\tau(k+1)-1}$, where the $k$ is determined by the relation $k=\lfloor l/\tau\rfloor$. This means that the successful transmission probabilities for WCS during these $\tau$ consecutive time instants $t_{l}$ depend on the state of MAS at time instant $t_{\tau k}$, i.e., $\alpha(k)$. Then, the WCS and MAS coupling is modeled by the following state-dependent fading channel:
$\mathbb{P}\Big{\{}\lambda_{i}(l)=1|\alpha(k)=\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}=\bar{\lambda}_{i}(\alpha),$ (5)
where $\lambda_{i}(l)=1$ indicates the successful transmission of link $i$ at time instant $t_{l}$.
Let $\eta_{i}(l)=1$ denote the successful packet decoding of link $i$. As shown in Fig. 2, with wireless communication policy (2), it holds
$\bar{\lambda}_{i}(\alpha)\!=\!\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)\mathbb{P}\{\eta_{i}(l)\!=\!1|\alpha(k)\!=\!\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\},$
where $\bar{\gamma}_{i}(c,\alpha)=\mathbb{P}\{\gamma_{i}(l)=c|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$ characterizes the probability distribution for the local channel state, and can be easily obtained by measuring the local fading amplitude for each MAS state by the statistical methods in [Agrawal2014, Kashiwagi2010]. Taking into account the interference among transmission links by adopting the collision channel model [gatsis2018], the last term in $\bar{\lambda}_{i}(\alpha)$ is equal to $\mathbb{P}\{\hbox{SNR}_{i}(l)\geq\eta_{i}^{0}|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor,\varphi_{i}(l)=1,\varphi_{j}(j)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(l)=0|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$, where $\hbox{SNR}_{i}(l)$ and $\eta_{i}^{0}$ denote the signal-to-noise ratio and the specified decodability threshold, respectively.
Figure 2: State-dependent fading channel (5).
Eq. (5) presents a general state-dependent channel model. In this model, the probability of a successful transmission for WCS (3) depends on the state of MAS (4). In addition, as shown in Fig. 2, Eq. (5) takes into account wireless communication policies, the interference among transmission links, and different time steps for the WCS and MAS, which cannot be captured by the existing channel models in [hu2022, hu2019].
@@ 2.6 Problem formulation
Considering the state-dependent fading channel model (5), given an initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ for MAS (4) such that $\alpha(k)\in\mathcal{C}_{\alpha}$, $k=0,1,\cdots$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
@@
The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}$ and a positive integer $L_{0}$ such that the inequalities
$\mathbb{E}\Big{[}V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\rho_{i}V_{i}(x_{i}(l))+Tr(Q_{i}\Xi_{i}),\forall x_{i}(l)\in\mathbb{R}^{n_{i}}$ (6)
hold for any $l\geq L_{0}$, any $x_{i}(l)\in\mathbb{R}^{n_{i}}$ and any $i\in\{1,\cdots,q\}$, where $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
In (6), $\rho_{i}$ represents the desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation. Similar kinds of performance requirements as in (6) have been adopted for WCSs [gatsis2015, gatsis2018]. Different from them, Definition 1 explicitly explores the effect of mobile agents’ movements on the successful wireless transmission of the WCS. In addition, it allows adopting different time steps to capture different time scales between the wireless transmission of control systems and the movements of mobile agents from cell to cell.
@@
$J_{{\bf{u}}}=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\Big{\{}\sum\limits_{l=0}^{K-1}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\Big{\{}\varphi_{i}(l)=1|\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda\sum\limits_{k=0}^{\lfloor(K-1)/\tau\rfloor}g(\alpha(k),u(k))\Big{\}}$ (7)
reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\mu_{i}$ is the transmit power for sensor $i$, $\sum_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS, and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
$P_{1}:$ $\min_{{\bf{u}}}J_{{\bf{u}}}$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(0)=\alpha_{0},$
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k)\in\mathcal{C}_{\alpha},u(k)\in\mathcal{C}_{u}(\alpha(k)),k=0,1,\cdots.$ (8)
3 Optimal wireless control as optimal constrained set stabilization
In the optimal control problem $P_{1}$, there are heterogeneous system dynamics constraints consisting of continuous component (3) and logical component (4) with different time steps. In addition, control performance (6) of the continuous system is constrained by the logical one. The disparity in analyzing these distinct systems makes the resolution of the problem difficult. To address these challenges, we convert WCS’s performance requirements (6) together with safety constraints (8) into the constrained set stabilization of MAS (4) using the state-dependent fading channel (5). In this way, problem $P_{1}$ is converted to the optimal constrained set stabilization problem of the MAS.
For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ an admissible input sequence in the following.